OpenStudy (anonymous):

A motor scooter purchased for $1,000 depreciates at an annual rate of 15%. Write an exponential function, and graph the function. Use the graph to predict when the value will fall below $100.

OpenStudy (anonymous):

You must be able to start this one off for yourself. Can you show us what you have so far?

OpenStudy (anonymous):

1000(1.15)=100?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

can u help me :)

OpenStudy (anonymous):

ok ask you self . what is the value of motor scooter after 1 year?

OpenStudy (anonymous):

150$....

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so what the question is really saying is , the value of motor scooter goes down by 15%.

OpenStudy (anonymous):

100 = 1000e^(-0.15t)

OpenStudy (anonymous):

850$

OpenStudy (anonymous):

@SNSDYoona we dont give answers out like that.

OpenStudy (anonymous):

1000 = the starting money that u purchased it the basic formula for this exponential decay and growth functions is A= Ye^(kt)

OpenStudy (anonymous):

thts why i give him the answer first to see if he does any calculation wrong then explain to him if he doesnt get it

OpenStudy (anonymous):

thats why we don't give answers out. you need to explain the exponential formula. Not just stated it!!!

OpenStudy (anonymous):

A = the amount u left after a certain period of time Y = initial amount started with k = ur rate of decay/growth t= the time

OpenStudy (anonymous):

0.1=(-0.15t)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

have u learn natural logs?

OpenStudy (anonymous):

from 100 = 1000e^(-0.15t) u simplify it then u get down to 0.1= e^(-0.15t) then u take the natural logs off both side u get ln(0.1) = -0.15t then u solve for t ln(0.1)/0.15 = t

OpenStudy (anonymous):

i got 0.666

OpenStudy (anonymous):

u shud get 15.35

OpenStudy (anonymous):

use ur calculator.. did u type in correctly?

OpenStudy (anonymous):

i divide .1 by .15?

OpenStudy (anonymous):

nonono u use ln << ln = natural log

OpenStudy (anonymous):

look at ur calculator and search for ln

OpenStudy (anonymous):

press ln and then (0.1)

OpenStudy (anonymous):

then divide by (-0.15)

OpenStudy (anonymous):

im using google calculator i dont have 1

OpenStudy (anonymous):

15.3?

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

v(t) = 0.85(1000)t; the value will fall below $100 in about 18 yr. v(t) = 1000(0.85)t; the value will fall below $100 in about 18 yr. v(t) = 0.85(1000)t; the value will fall below $100 in about 14.2 yr. v(t) = 1000(0.85)t; the value will fall below $100 in about 14.2 yr.

OpenStudy (anonymous):

u still there

OpenStudy (anonymous):

hi

OpenStudy (liammcbay):

hi